VIBRATIONAL SPECTROSCOPY AND DYNAMICS OF THE HYDRAZOIC AND ISOTHIOCYANIC ACIDS IN PROTIC AND APROTIC SOLVENTS

\underline{C. HOUCHINS}, D. WEIDINGER^{a}, Code 6111, U.S. Naval Research Laboratory, Washington, DC 20375-5342; D. BROWN, U.S. Naval Academy, Chemistry Department, Annapolis, MD 21402; J. OWRUTSKY, Code 6111, U.S. Naval Research Laboratory, Washington, DC 20375-5342.

In order to investigate the effects of solute charge as well as solvent-isotope effects on solute vibrational spectra and dynamics, infrared pump-probe studies have been carried out to determine vibrational energy relaxation (VER) times for the CN and antisymmetric NNN stretching bands for the pseudohalide acids, XNCS and XN_{3} (X=H, D), in protic and aprotic solvents to compare with the well studied azide and thiocyanide anions. The studies indicate that the deuterium effects on the frequency shifts for HN_{3} and relaxation times for both HNCS and HN_{3} resemble the solvent isotope effects for azide and thiocyanide in water. While it is expected that VER times will be shorter for charged than neutral solutes, this is not observed in all cases.

aCH and DW acknowledge the Naval Research Laboratory - National Research Council Research Associateship.